# Policy Learning with Confidence

### v2

@@ 1. Introduction
Notes: Hollow squares represent MVPF estimates and their standard errors for 14 programs collected by the Policy Impact Library. The red curve shows the decision frontier for budget allocation across these programs. The solid square and circle mark allocations selected by EWM and PoLeCe, respectively. For programs targeting age $>$ 25, the programs included in the PoLeCe allocation with large shares, listed in order of their shares, are: Housing Vouchers in Chicago (0.65 $\pm$ 0.03), OHIE, Single Adults (1.16 $\pm$ 0.04), where each value represents the point estimate ($\pm$ standard error). The program selected by EWM is Holistic Wrap-around Services Can Improve Employment Rates (1.39 $\pm$ 1.11). For programs targeting age $\leq$ 25, the programs included in the PoLeCe allocation are: Head Start (1.84 $\pm$ 0.47) and Wisconsin Scholars Grant (1.43 $\pm$ 0.38). The program selected by EWM is Head Start (1.84 $\pm$ 0.47).
This tradeoff is illustrated in Figure 1 using the MVPF estimates from the Policy Impacts Library, which comprises 14 US programs after we focus on estimates based on randomized control trials (RCTs) with finite standard errors. To mimic a DM who applies constant welfare weights within certain age groups but not necessarily across them due to distributional concerns, we use the Policy Impacts Library classification of programs based on the intended beneficiaries’ age. We then consider budget allocation separately for two distinct age groups. Six of these programs target beneficiaries below age 25, covering early childhood education programs, college financial aid, and job training. [^5] As explained in the Policy Impacts Library, alternative specifications have led to larger MVPF estimates for early childhood education programs depending on how lifetime earnings are forecasted and whether one accounts for the transfer value of preschool subsidies to parents. Eight of these programs target beneficiaries above age 25, covering federal social assistance programs, health insurance, and housing vouchers. In both cases, the MVPF of the programs selected by the EWM rule are not estimated as precisely as the welfare delivered by diversifying the budget across several programs whose MVPFs are estimated with significantly higher precision. If a DM values both welfare and precision, how should these goals be balanced?
We propose Policy Learning with Confidence, or PoLeCe, among the class of risk-aware policy rules that lie on the efficient decision frontier. Like all risk-aware rules, PoLeCe uses observable measures of statistical variation to adjust policy choice relative to a simple plug-in rule by penalizing the observed performance of each policy by an amount proportional to its standard error. Specifically PoLeCe is a risk-aware decision rule that uses a data-dependent penalization factor so that the resulting objective value automatically provides the lower bound of a one-sided confidence interval for the welfare achieved by the selected rule. Thus, the rule delivers a reporting guarantee, ensuring with high confidence that the actual welfare exceeds a lower threshold; no adjustment for post-selection inference is required. Importantly, policies that allocate fractional shares to different treatments or programs can be allowed, producing a richer frontier than that available from singleton allocations such as the individual programs indicated by hollow squares in Figure 1.
In the above example, all that was required to map out the efficient decision frontier, and indeed all that is needed to determine the optimal allocation following our proposal, are point estimates for the welfare of each available policy and their joint variance-covariance. [^6] Our proposed PoLeCe rule uses the sample correlation of the estimates to calibrate the precision penalty to achieve the aforementioned reporting guarantee. The general class of risk-aware decision rules described in Section 2.1 allows for different choices for the penalization factor, which corresponds to the green slope highlighted in Figure 1. Other risk-aware rules may thus only require standard errors rather than estimates of the entire correlation structure, depending on their choice of penalization factor. This can be obtained in applications that feature a variety of different models and sampling processes. An important special case to which our framework applies is the analysis of optimal treatment assignment using data on individual-specific treatments and allocations, such as from an RCT. For example, in a real-world setting of treatment assignment, while GiveWell uses EWM based on RCT results, they have recently highlighted the importance of accounting for estimation uncertainty in decision-making for the sake of “transparency” (Salisbury, 2024). Our proposal offers a practical solution to this concern.
The analysis of statistical treatment rules in such settings has received renewed attention in the recent literature, starting with the pioneering work of Manski (2004), who considers treatment choice on the basis of minimax (expected) regret. Follow-up works include a decision theoretic framework introduced by Dehejia (2005), finite-sample bounds considered by Stoye (2009), and an asymptotic framework introduced by Hirano and Porter (2009, 2020). Hypothesis testing is a conventional approach to decision-making under uncertainty. As argued by Tetenov (2012), it can be interpreted as a DM having asymmetric aversion to Type I and Type II regret.
Kitagawa and Tetenov (2018) proposes empirical welfare maximization (EWM) rules for experimental data, which is an “as-if” optimization rule that selects a policy with the highest estimated welfare (empirical welfare), thus constituting a plug-in rule for treatment effect models with experimental data. Kitagawa and Tetenov (2018) proves the optimality of EWM rules in the sense that as the sample size increases expected regret converges to zero at the minimax rate. Athey and Wager (2021) and Demirer et al. (2019) propose doubly-robust estimation of average welfare, in which case this result obtains even with quasi-experimental data. In the context of this literature, PoLeCe offers a modification of EWM that incorporates estimation risk into the decision rule. The EWM rule corresponds to the PoLeCe rule with penalization factor of zero on the standard error, putting no weight on estimation risk. Inference approaches for optimal treatment assignment and/or welfare complementary to EWM, but which do not propose alternative assignment rules to account for uncertainty, include those of Armstrong and Shen (2023), Ponomarev and Semenova (2024), and Rai (2019), as well as the online supplement of the original paper by Kitagawa and Tetenov (2018). As an extension of EWM, Mbakop and Tabord-Meehan (2021) proposes a penalized welfare maximization rule that penalizes for the complexity of the policy, and establishes an oracle property for model selection.
Andrews and Chen (2025) proposes the use of certified decisions, which comprise the pairing of a decision rule with an accompanying inferential guarantee. Their results provide a decision theoretic justification for the PoLeCe rule, showing that it belongs to an essentially complete class of decision rules for a DM who requires a coverage guarantee, and therefore dominates the pairing of EWM with the studentized projection approach in Andrews et al. (2024).
PoLeCe also contributes to a recent literature featuring some alternative proposals to directly account for estimation uncertainty in the decision rule. In settings with a budget constraint, Sun (2021) proposes a rule that penalizes policies with large uncertainty in meeting the constraint, and Moon (2025) proposes using empirical Bayes estimates, leveraging shared information across related policies. Kitagawa et al. (2022) focuses on the variability of regret and shows that the EWM rule is dominated by fractional rules when the objective function is the sum of squared expected regret and the variance of regret. If the estimated welfare is independent across policies, then one may appeal to the literature on optimal selection such as Gupta and Miescke (1986), referenced above. However, in the context of policy learning, since the groups assigned by different policies might not be mutually exclusive, the estimated welfare associated with different policies may not be independent. For instance, when treatment allocation is determined by an income threshold, the welfare estimates for groups falling below different thresholds will be correlated by construction.
The paper proceeds as follows. Section 2 sets out our approach to risk-aware policy choice that selects policies from the estimated efficient decision frontier. Section 2.1 motivates the approach using selection of policies for treatment choice and investment allocation in public programs as key illustrative examples. Section 2.2 develops favorable regret and lower confidence guarantees for the general class of risk-aware policy rules, and Section 2.3 formally sets out the proposed PoLeCe rule as the risk-aware rule that maximizes a lower confidence bound on maximal welfare. Section 2.4 provides evidence from Monte Carlo experiments calibrated to empirical applications. Section 2.5 provides alternative motivations for risk-aware decisions. Section 2.6 discusses refinements and further theoretical points. Section 3 provides four empirical applications that demonstrate the application of PoLeCe, and Section 4 concludes.
2. Risk-Aware Policy Choices
2.1. Motivation and Risk-Aware Policy Choice
We begin with the following key problem as motivation.
Problem of Choosing Better Treatment Policies.
@@ 2.1. Motivation and Risk-Aware Policy Choice
Risk-Aware Policy Choices
Different treatment policies $\pi$ can have markedly different welfare and estimation risk profiles, and some may be much riskier than others. For example, policies that “turn on” treatments with near-zero propensities may be especially risky, since
$s^{2}(\pi)\propto\mathrm{E}[H^{2}(\pi)]/n$
can be large in such cases. What should a decision maker do?
The decision maker has to choose an optimal policy in $\Pi$ based on data $\widehat{V}(\pi)$ that is (approximately) Gaussian. We argue that an optimal risk-aware decision can be achieved by maximizing the empirical welfare offset by the estimation risk times a critical value:
$\widehat{\pi}_{\text{RW}}\;\in\;\arg\max\Bigl{\{}\widehat{V}(\pi)-k\,\widehat{s}(\pi):\pi\in\Pi\Bigr{\}},$ (4)
where $k$ is a critical value and $\widehat{s}(\pi)$ is a consistent estimator of the estimation risk $s(\pi)$. By varying $k$, we trace out the efficient decision frontier of treatment policies, as illustrated by Figure 1 in the introduction. Each point on the frontier corresponds to a particular $k>0$. We provide fast algorithms in Appendix B to efficiently compute the frontier for any application.
Our leading proposal to choose $k$ is to meet certain reporting guarantees with high confidence, thereby conducting “policy learning with confidence” (PoLeCe). [^9] Other choices of $k$ are of course possible and can be rationalized as optimal under risk aversion. We discuss those later in Section 2.5. The resulting choice generally differs from the empirical welfare maximizer (EWM):
$\widehat{\pi}_{\text{EWM}}\;\in\;\arg\max\{\widehat{V}(\pi):\pi\in\Pi\}.$ (5)
The EWM rule arises when $k\searrow 0$. We provide a comparative analysis of these two rules in what follows. Before we present details of our PoLeCe proposal, we discuss another class of applications.
The Problem of Choosing Investment Allocations in Public Programs
A public or private agency is considering programs $\{1,...,J\}$ and must decide on investment shares
$\pi=(\pi_{1},...,\pi_{J}),$
subject to
@@
which is the unit simplex intersected with a rectangular set of constraints. These additional constraints may reflect diversity or other requirements that allocations must satisfy. The welfare of the allocation $\pi$ is
$V(\pi)=\pi^{\prime}R,$
@@
so that the estimated welfare of the allocations are also approximately Gaussian:
@@
where $\{Z_{\pi}\}_{\pi\in\Pi}$ is a Gaussian vector with standard normal marginal distributions.
Hence, for each fund allocation $\pi$, the agency is given $\widehat{V}(\pi)$ along with its associated estimation risk $s(\pi)$. What should the agency do? Our proposal is again to use the risk-aware rule (4). If estimation risk is ignored, one might rely on the empirical welfare maximizer, but in the absence of additional constraints, that rule would simply allocate all funds to the program with the highest estimated return. This is generally unappealing on intuitive grounds. In contrast, the risk-aware approach (4) would spread out funds in an “efficient” portfolio of programs, similar to Markowitz portfolio allocation in finance. However, the motivation and formulation of our approach are distinct from the Markowitz model.
2.2. Regret Properties and Guarantees of Any Risk-Aware Policy
We now provide regret bounds for any risk-aware decision rule $\widehat{\pi}_{\text{RW}}(\widehat{k})$ that solves the risk-adjusted empirical welfare problem:
We focus on the case where $\Pi$ has finite cardinality $p=p_{n}$, which can be very large, potentially much larger than $n$. (A theoretical treatment for infinite $\Pi$ follows by suitable discretization, as we comment on later.)
Scaled Estimation Error and Gaussian Approximation. Define
$\widehat{Z}_{\pi}\;:=\;\frac{\widehat{V}(\pi)-V(\pi)}{\widehat{s}(\pi)}$
to be the scaled estimation error process. We use the following Gaussian approximation condition, denoted by (G). Let $\mathcal{A}$ represent the collection of rectangular sets in $\mathbb{R}^{p}$.
  • (G)
    $\Bigl{\lvert}\Pr\!\bigl{(}(\widehat{Z}_{\pi})_{\pi\in\Pi}\in A\bigr{)}\;-\;\Pr\!\bigl{(}(Z_{\pi})_{\pi\in\Pi}\in A\bigr{)}\Bigr{\rvert}\;\leq\;r_{n},\quad\text{ for all }A\subset\mathcal{A}.$ (G)
These conditions are known to hold under mild assumptions. [^11] See Chernozhukov et al. (2022, 2023a) for sharp forms of Gaussian approximations, and Chernozhukov et al. (2023b) for a review. Additional references, e.g. Belloni et al. (2018); Quintas-Martinez (2022) discuss Gaussian approximations for many or a continuum of target parameters, including those learned via debiased machine learning. Furthermore,  (G) implicitly requires consistency of risk estimates: $\max_{\pi\in\Pi}\lvert\widehat{s}(\pi)/s(\pi)-1\rvert\to_{P}0$.
denote the $(1-\beta)$-quantile of the estimation error over $K$ under the Gaussian approximation. Under (G), with probability at least $1-\beta-r_{n}$, $\max_{\pi\in K}\widehat{Z}_{\pi}\leq q_{1-\beta,K}$. Let $V_{\max}=\sup_{\pi\in\Pi}V(\pi)$ be the maximal true welfare and $\Pi_{0}=\{\pi\in\Pi:V(\pi)=V_{\max}\}$ be the set of policies with maximal true welfare, and denote its cardinality by $p_{0}=|\Pi_{0}|$. Define
Proposition 1 (Regret Bounds for Risk-Aware Decisions).
Under condition (G), the regret of any RW policy selection rule $\widehat{\pi}_{\text{RW}}(\widehat{k})$ is bounded with probability at least $1-2\beta-2r_{n}$ as:
$V_{\max}-V(\widehat{\pi}_{\text{RW}}(\widehat{k}))\;\;\leq\;\;\frac{\underline{\sigma}_{\Pi_{0}}}{\sqrt{n}}\Bigl{\{}q_{1-\beta,\Pi_{0}}+\widehat{k}\Bigr{\}}\;+\;\frac{\bar{\sigma}_{\Pi}}{\sqrt{n}}\Bigl{\{}\bigl{(}q_{1-\beta,\Pi}-\widehat{k}\bigr{)}_{+}\Bigr{\}}.$
As $\beta$ decreases, the probability that the regret bound holds increases, since the failure probability is at most $2\beta+2r_{n}$. However, this comes at the cost of a looser bound because the quantile $q_{1-\beta,\Pi}$ increases as $\beta$ gets smaller. This reflects a basic trade-off: a smaller $\beta$ provides a more conservative bound that holds with higher probability, while a larger $\beta$ results in a tighter bound that is less certain to be valid. This behavior is analogous to confidence intervals, where greater confidence requires a wider interval.
Proposition 1 provides dimension-free bounds. To obtain simpler dimension-based bounds, observe that by standard concentration inequalities, [^12] We can also employ Talagrand’s style upper bounds, but to make our points it suffices to use these simple concentration bounds.
We now see that the broad class of RW-rules with limited risk aversion has regret bounded by
$\bar{\sigma}_{\Pi}\,\sqrt{\frac{\log p}{n}},$
as $p\to\infty$ (and $n\to\infty$). This class includes the empirical welfare maximization rule ($k=0$) and matches known optimal minimax rates when available (see Athey and Wager (2021)). Additionally, choosing $\widehat{k}>0$ sufficiently large can yield a regret bound of
$\underline{\sigma}_{\Pi_{0}}\,\sqrt{\frac{\log p}{n}},$
$\underline{\sigma}_{\Pi_{0}}<\bar{\sigma}_{\Pi}$
if there is significant heterogeneity in estimation risks across policies. The empirical risk minimization literature has also used penalties on estimation risk to improve error bounds, albeit in different contexts; see, for example, Maurer and Pontil (2009); Swaminathan and Joachims (2015); Foster and Syrgkanis (2023).
Moving Toward a Data-Driven $\widehat{k}$. These ideas naturally lead to the next section, where we propose a main policy rule that carefully constructs a data-driven $\widehat{k}\approx q_{1-\alpha,\Pi}$ to achieve small regret and provide additional reporting guarantees.
Reporting Guarantees of Risk-Aware Decisions
In practice, to obtain reporting guarantees for RW decisions, one needs to approximate the quantiles $q_{1-\alpha,\Pi}$. We do so via the bootstrap, where we set:
where $\widehat{Z}_{\pi}^{*}:=N\bigl{(}0,\widehat{C}\bigr{)}$, and $\widehat{C}$ is a consistent estimator of the covariance matrix $C=\mathrm{Cov}((Z_{\pi})_{\pi\in\Pi})$. Denote by $\Pr^{*}$ the bootstrap-induced probability measure computed conditional on $\widehat{C}$. This construction relies on the following condition:
  • (B)
    With probability at least $1-\delta_{n}$ (where $\delta_{n}\searrow 0$),
    $\Bigl{\lvert}{\Pr}^{*}\!\bigl{(}(\widehat{Z}_{\pi}^{*})_{\pi\in\Pi}\in A\bigr{)}\;-\;\Pr\!\bigl{(}(Z_{\pi})_{\pi\in\Pi}\in A\bigr{)}\Bigr{\rvert}\;\leq\;r_{n},\quad\text{for all }A\subset\mathcal{A}.$ (B)
Like (G), condition (B) is satisfied under mild assumptions even when $p$ is much larger than $n$, and has been verified for a variety of estimation methods, including debiased machine learning. [^13] See Chernozhukov et al. (2022, 2023a) and Chernozhukov et al. (2023b) for more details, along with other references on “approximate means” settings such as debiased ML (Belloni et al. (2018); Quintas-Martinez (2022)).
Proposition 2 (LCB Guarantees for RW Decisions).
2.3. The PoLeCe Rule and Its Properties
$\widehat{\pi}_{\mathrm{PoLeCe}}\;\;\in\;\;\arg\max_{\pi\in\Pi}LV_{1-\alpha}(\pi).$ (7)
Notably, PoLeCe is an RW decision that uses
$\widehat{k}=\widehat{q}_{1-\alpha,\Pi}$
(the bootstrap quantile) as the penalty for estimation risk. Furthermore, the maximized LCB
serves as a natural performance guarantee.
Remark 1 (PoLeCe as Minimizer of the UCB on Regret).
Proposition 3 (Reporting Guarantees and Regret Bounds for PoLeCe).
With probability at least $1-\alpha-r_{n}^{\prime}$,
$V_{\max}\;\;\geq\;\;V(\widehat{\pi}_{\mathrm{PoLeCe}})\;\;\geq\;\;LV_{1-\alpha,\Pi}.$
Moreover, with probability at least $1-\alpha-\beta-r_{n}-r_{n}^{\prime}-\delta_{n}$, the regret of $\widehat{\pi}_{\mathrm{PoLeCe}}$ satisfies
The first result in Proposition 3 shows that $LV_{1-\alpha,\Pi}$ provides a valid high-confidence lower bound on the true policy welfare of the PoLeCe choice. The second result sharpens the general regret bounds presented in Proposition 1. As noted following Proposition 1, the parameter $\beta$ governs the trade-off between failure probability and bound tightness, while the constant $\alpha$, chosen by the researcher, determines the confidence level of the lower bound $LV_{1-\alpha,\Pi}$. The refinement in Proposition 3 depends on the relationship between $\alpha$ and $\beta$. When $\alpha=\beta$, Propositions 1 and 3 yield qualitatively identical guarantees. When $\alpha<\beta$, the bounds in both propositions remain qualitatively identical, but the bound in Proposition 3 holds with higher probability. When $\alpha>\beta$, the bound in Proposition 3 is tighter, but it may hold with lower probability. For a large $\alpha$, the bound may fail to hold; in such cases, the general bound from Proposition 1 remains valid and informative, as it applies to all risk-aware decision rules.
Another key observation is that the leading constant in the regret bound is governed by the least risky among the best policies, and can therefore be significantly smaller than that for the EWM rule in settings with substantial risk heterogeneity. This theoretical advantage is corroborated by the computational experiments we present below.
2.4. Computational Experiments

### v3

@@ Appendix C Exhibits
Notes: Squares represent MVPF estimates and their standard errors for 14 programs collected by the Policy Impacts Library. The red curve shows the decision frontier for budget allocation across these programs. The solid square and circle mark allocations selected by EWM and PoLeCe, respectively. For programs targeting age >> 25, the programs included in the PoLeCe allocation, listed in order of their shares, are: OHIE, Single Adults: 1.16 (0.04), Job Training Partnership Act, Adults: 1.38 (0.40), where each value represents the point estimate (standard error). The program selected by EWM is Holistic Wrap-around Services Can Improve Employment Rates: 1.39 (1.11). For programs targeting age $\leq$ 25, the programs included in the PoLeCe allocation are: Head Start: 1.84 (0.47) and Wisconsin Scholars Grant: 1.43 (0.38). The program selected by EWM is Head Start: 1.84 (0.47).
Table 1. Results for Investment Allocations to Adult Programs (Age >> 25)
EWM PoLeCe
Panel A. Outcome: MVPF
[…]
Panel B. Selected Policy
Padua Pilot: 0.18 (0.23) 0 0.00
Housing Vouchers in Chicago: 0.65 (0.02) 0 0.00
Oregon Health Insurance Experiment (Provided to Single Adults): 1.16 (0.04) 0 0.97
Paycheck Plus: EITC to Adults without Dependents: 1.00 (0.09) 0 0.00
Work Advance: 0.78 (0.28) 0 0.00
Job Training Partnership Act, Adults: 1.38 (0.40) 0 0.03
EWM PoLeCe
Panel A. Outcome: MVPF
[…]
Panel B. Selected Policy
Head Start Impact Study: 1.84 (0.47) 1 0.52
JobStart: 0.20 (0.11) 0 0
Year Up: 0.43 (0.03) 0 0
Job Corps: 0.15 (0.22) 0 0
Job Training Partnership Act, Youth: -0.23 (0.81) 0 0
Online Appendix to “Policy Learning with Confidence”
Victor Chernozhukov, Sokbae Lee, Adam M. Rosen and Liyang Sun
Appendix D Details for Gaussian Bootstrap Implementation
The following Lemma provides formalities for implementing the Gaussian bootstrap procedure described in Appendix B.
Lemma 2 (Existence, Uniqueness, and Differentiability of the Root).
Then:
  1. (1)
    For each fixed $t\geq 0$, the optimization problem
  2. (2)
    There exists a unique $t^{*}\geq 0$ such that $f(t^{*})=0$.
  3. (3)
    The derivative of $f$ at $t$ is given by
    $f^{\prime}(t)=-\,\sqrt{\,\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)\,}$,
Implications for Gaussian Bootstrap
The reformulation of computing $\bar{Z}^{*}$ as a root-finding problem over a family of second-order cone programs significantly enhances computational efficiency. Each evaluation of $f(t)$ involves solving a concave SOC optimization problem, which is well-supported by modern optimization solvers. The uniqueness of the root $t^{*}$ guarantees that iterative root-finding algorithms, such as the Newton-Raphson method or bisection, will converge reliably to the desired solution. Moreover, the explicit expression for the derivative $f^{\prime}(t)$ facilitates the use of derivative-based optimization methods, further streamlining the computational process.
Proof of Lemma 2.
1. Second-Order Conic Reformulation. Fix any $t\geq 0$. Consider
Introduce an auxiliary variable $s$ such that
The optimization problem becomes
$\max_{\pi,s}$ $\pi^{\top}z+t\,s$
subject to $\|\widehat{\Omega}^{1/2}\,\pi\|_{2}\leq-s,$
$s\leq 0,$
$\pi\in\Pi.$
2. Existence and Uniqueness of the Root. Existence. Observe that
$f(0)=\max_{\pi\in\Pi}\pi^{\top}z,$
which is finite since $\Pi$ is bounded and closed. As $t\to\infty$, the term $t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}$ dominates for every $\pi\in\Pi$, leading to $f(t)\to$-$\infty$. Since $f(t)$ is continuous (as the pointwise maximum of continuous functions over a compact set), the Intermediate Value Theorem guarantees the existence of some $t^{*}\geq 0$ such that $f(t^{*})=0$.
Uniqueness. From part 3, we have
$f^{\prime}(t)=-\,\sqrt{\pi^{*}(t)^{\top}\widehat{\Omega}\,\pi^{*}(t)}<0,$
3. Derivative of $f(t)$. Define
$F(\pi,t)=\pi^{\top}z-t\,\sqrt{\pi^{\top}\widehat{\Omega}\,\pi}.$
Then
$f(t)=\max_{\pi\in\Pi}F(\pi,t).$
By the envelope theorem, the derivative of $f(t)$ with respect to $t$ is
Appendix E Alternative Motivations for Risk-Aware Decisions
We consider a decision maker (DM) who holds posterior beliefs $B$ about policy welfare $\{V(\pi)\}$ given by
$V(\pi)\;\sim\;\widehat{V}(\pi)\;+\;\widehat{s}(\pi)\,Z_{\pi},$ (7)
$\max_{\pi\in\Pi}\,\mathrm{E}_{B}\bigl[U\bigl(V(\pi)\bigr)\bigr].$
Proposition 4 (Posterior Expected Utility Maximization).
  1. (1)
    for some constant $k_{UB}$ that depends on $U$ and $B$.
  2. (2)
$\mathrm{E}[U(V(\pi))]\;=\;1\;-\;\exp\Bigl(-\lambda\,\widehat{V}(\pi)\;+\;\tfrac{1}{2}\,\lambda^{2}\,\widehat{s}^{2}(\pi)\Bigr).$
Consider the loss function
$V_{\max}-V(\pi)=\max_{\bar{\pi}\in\Pi}V(\bar{\pi})-V(\pi)$
$\min_{\pi\in\Pi}\mathrm{E}_{B}\bigl[L(V(\pi))\bigr].$
Proposition 5 (Best Posterior Regret-Risk Aware Decision).
$\max_{\pi\in\Pi}\,\Bigl\{\,\widehat{V}(\pi)\;-\;k_{R}\,\widehat{s}(\pi)\Bigr\}.$
Appendix F Proofs of Propositions 4 and 5
Proof of Proposition 4
Claim (1). Let $v=\widehat{V}(\pi)$ and $s=\widehat{s}(\pi)$. Then
$\frac{\partial}{\partial v}f(v,s)\;=\;\mathrm{E}[U^{\prime}(v+sZ)]\;>\;0,\quad\frac{\partial}{\partial s}f(v,s)\;=\;\mathrm{E}\bigl[U^{\prime}(v+sZ)\,Z\bigr]\;<\;0,$
because $U^{\prime}(\cdot)>0$ and $U^{\prime}(v+sZ)$ is comonotonic with $Z$ (given that $U^{\prime\prime}$ exists by Alexandrov’s theorem almost everywhere and $U^{\prime\prime}<0$ by strict concavity). The second inequality follows from Chebyshev’s association inequality, and the first holds because the expectation of a strictly positive random variable is positive.
[…]
This maximization can be written in a Lagrangian form:
[…]
Proof of Proposition 5.
Observing that
and
the program becomes
Appendix G Efficient Decision Frontier Computation
In this appendix, we provide pseudocode of the algorithm to compute the efficient decision frontier. The code below builds a monotonically nondecreasing, concave frontier from the input of points.
Algorithm 1 Build a Monotonically Nondecreasing, Concave Frontier
1:$points=(Risk,Value)$ such that $Risk$ and $Value$ have the same length
2:The set of points that form the efficient frontier
3:procedure FrontierEstimation(pointspoints)
4:  Create data structure from input
5:  Sort by ascending RiskRisk, tie-break by descending ValueValue
6:  $frontier\leftarrow$ empty list
7:  for $i\leftarrow 1$ to the number of rows of pointspoints do
8:   $p\leftarrow\text{row $i$ of $points$}$ $\triangleright$ extract row $i$ from pointspoints
9:   if frontierfrontier is not empty and $p.Value<frontier[\text{last}].Value$ then
11:   end if
12:   while there are at least two elements in frontierfrontier do
13:     $p2\leftarrow\text{last frontier point}$
14:     $p1\leftarrow\text{second to last frontier point}$
16:      remove $p2$ from frontierfrontier $\triangleright$ ensure concavity
17:     else
18:      break
19:     end if
20:   end while
21:   append $p$ to frontierfrontier
22:  end for
23:  return frontierfrontier $\triangleright$ the set of frontier points
24:end procedure
For the investment allocation problem, it may be infeasible to discretize $\Pi$ and apply the above algorithm to plot the efficiency frontier. Instead, we start with a fine grid for $k$ centered at $\hat{q}_{0.95,\Pi},$ and directly trace out the efficiency decision frontier by solving the corresponding optimization problem as in  (2). Appendix B provides a computationally efficient way to solve such optimization and therefore to plot the efficiency frontier.
Appendix H Treatment Choice
This appendix considers the problem of selecting treatment choice policies. Section H.1 describes the treatment choice problem, and Section H.2 then demonstrates application of PoLeCe in three empirical settings using data from RCTs.
H.1. Problem of Choosing Better Treatment Policies.
@@ Appendix H Treatment Choice
$\{\widehat{V}(\pi)\}\;\overset{a}{\sim}\;\{V(\pi)+s(\pi)\,Z_{\pi}\},\quad Z_{\pi}\sim N(0,1),\quad s^{2}(\pi)=\mathrm{Var}(\psi_{\pi})/n,$
where $\{Z_{\pi}\}_{\pi\in\Pi}$ is a centered Gaussian process, with each component having standard normal distribution; and $\overset{a}{\sim}$ means “approximately distributed as” in the sense of Condition (G) in Section 2.2.
H.2. Empirical Applications to Treatment Choice
In this section we demonstrate application of PoLeCe to the problem of choosing better treatment policies. We use RCT data from Banerjee et al. (2025), Dupas and Robinson (2013), and Schilbach (2019) to illustrate the usefulness of our method and to show how the results can be incorporated into research papers that analyze RCTs. These papers focus on estimating the added value relative to control (as defined in (10)). In contrast, we focus on each policy’s value to fully account for the estimation error for estimating the average outcome under control. While these papers originally used regression adjustment, we rely on efficient influence functions (doubly robust scores via (11)) to obtain welfare estimates for potential precision gains.
H.2.1. Immunization Nudges
(a) Measles shots outcome
(b) Shots per dollar outcome
Figure 2. Welfare estimates against precision
Table 3. Results for Immunization Nudges
EWM PoLeCe Control
Panel A. Outcome: Measles shots
@@ H.2.1. Immunization Nudges
Panel B. Outcome: Shots per dollar
@@
Panel C. Selected Policy:
Measles shots trusted info hub, info hub, no info,
high slope low slope no incentive
high SMS low SMS no reminder
Shots per dollar trusted info hub, trusted info hub, no info,
no incentive no incentive no incentive
no reminder no reminder no reminder
H.2.2. Informal Savings Technologies
Table 4. Results for Informal Savings Technologies
EWM PoLeCe Control
Panel A. Outcome: Health Investment (in Ksh)
@@ H.2.2. Informal Savings Technologies
[…]
Panel B. Selected Policy:
(Female, Married) =(0,0)=(0,0) Health Pot Health Pot Control
(Female, Married) =(0,1)=(0,1) Health Pot Health Pot Control
(Female, Married) =(1,0)=(1,0) Health Pot Safe Box Control
(Female, Married) =(1,1)=(1,1) Safe Box Health Pot Control
There are $|\Pi|=625=5^{4}$ possible treatment policies all together because there are five treatment values and four possible values in the support of the two covariates. Figure 3(a) plots all welfare estimates relative to their precision. The highest welfare is observed below 900 Ksh, but it is accompanied by the largest standard error. Therefore, it is advisable to account for the precision of these estimates when selecting an optimal policy. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 4 summarizes the empirical results. In Panel A, the maximized welfare by EWM is 890 Ksh, which corresponds to the highest welfare in Figure 3(a). The adjusted estimate of maximum welfare $\widehat{V}(\widehat{\pi}_{\mathrm{PoLeCe}})$ at $\alpha=0.05$ is 587, which is much less than no-precision-corrected $\widehat{V}(\widehat{\pi}_{\text{EWM}})$ but substantially greater than $\widehat{V}(\widehat{\pi}_{\text{EWM}})-\widehat{q}_{0.95,\Pi}\widehat{s}(\widehat{\pi}_{\text{EWM}})$. Furthermore, the chosen optimal policies differ for females. EWM selects Health Pot for unmarried women and Safe Box for married women. In contrast, after accounting for standard errors, PoLeCe recommends Safe Box for unmarried women and Health Pot for married women.
(a) Health investment (in Ksh)
(b) Labor supply
Figure 3. Welfare estimates against precision
H.2.3. Alcohol and Self-Control
As an illustration, we focus on labor supply measured by the fraction of days individuals worked from day 5 (the first day of sobriety incentives) through day 19 (the last day of sobriety incentives). For simplicity, we choose two baseline covariates: Baseline sober =1=1 if the baseline fraction sober is 1 and Baseline sober =0=0 otherwise; Owns rickshaw =1=1 if a driver own a rickshaw and Owns rickshaw =0=0 otherwise. After removing observations with missing values, there are 222 individuals. Given complete randomization, we calculate doubly robust scores based on the two baseline covariates.
Table 5. Results for Self-control
EWM PoLeCe Control
Panel A. Outcome: Labor Supply
[…]
Panel B. Selected Policy:
(Baseline sober, Owns rickshaw) =(0,0)=(0,0) Control Control Control
(Baseline sober, Owns rickshaw) =(0,1)=(0,1) Control Incentive Control
(Baseline sober, Owns rickshaw) =(1,0)=(1,0) Choice Choice Control
(Baseline sober, Owns rickshaw) =(1,1)=(1,1) Choice Choice Control
There are $|\Pi|=81=3^{4}$ possible treatment policies in total since there are three treatment levels and four possible values of the two covariates. Figure 3(b) plots the welfare estimates against their precision. Although the highest welfare appears with a relatively small standard error, this does not necessarily imply statistical significance. To implement PoLeCe at the level $\alpha=0.05$ and calculate $\widehat{q}_{0.95,\Pi}$, we apply the bootstrap procedure from Condition (B). Table 5 summarizes the empirical results. The EWM and PoLeCe estimates, along with their lower confidence bounds, are very similar and only slightly exceed those of the control group.
Appendix I Computational Experiments